Watch later Share Info Sample of 25 Fish Sample mean-X = 41 cm Sample SD 12 cm 10 20 30 40 50 60 Length (cm) RE VIDEOSLength (cm) SAMPLE MEAN MARGIN OF ERROR -> SAMPLE HEAN /- 2 X Standard erron of the mean SAMPLE MEAN/-2 X (SD n 11cm- 2 x (12Question 2 The margin of error in the video was 4.8cm. Using the critical value of standard normal, re-calculate the margin o
Answer #1

2. Margin of error using critical value of standard normal:

z*standard deviation/√n = 1.96*12/√25 = 4.704

So the answer is 4.7

3. Using t critical value, t score is t 0.025,24 = 2.064

So margin of error= 2.064*12/√25 = 4.9536

So rounding to the nearest tenth, the answer will be 5.0